# IR/UV mixing from higher-order interactions in a Scalar Field

### v1

@@ Abstract
In this work I present a concrete four-dimensional scalar field theory with a quasi-local, marginal (by power-counting) four-field interaction whose momentum-dependent structure generates a dominant $k^{8}$ contribution to the effective Hamiltonian of each Fourier mode at large wave number $k$. When the field is confined to a finite spacetime box of side $L$, the resulting dynamics render each high-$k$ mode an effective quartic oscillator rather than a harmonic one. A semiclassical estimate of the ground-state energy then enforces an emergent UV cutoff
@@
I discuss stability, the approximations used, and generalizations to other higher-order interaction structures. I then compute the resulting vacuum-energy density using dimensional regularization and propose a possible way to deduce the scaling of the effective equation of state with $L$. I then include qualitative comparison with the DESI indication of mildly time-varying dark energy. The aim is to provide a field-theoretic mechanism illustrating how quasi-local interactions can generate UV/IR mixing and dynamically modify the UV cutoff scale.
@@ I Introduction
$\rho_{\rm vac}^{\rm naive}\sim\frac{1}{2}\int^{\Lambda}\!\frac{d^{3}\bm{k}}{(2\pi)^{3}}\,\omega_{k}\;\sim\;\Lambda^{4}$ (1)
with $\omega_{k}=(k^{2}+m^{2})^{1/2}$. Choosing $\Lambda$ near the Planck scale overshoots the observed vacuum energy by about 120 orders of magnitude. Lorentz-invariant regularization schemes such as dimensional regularization preserve $p=-\rho$ for the vacuum state and significantly reduce the result, yet still leave a discrepancy Akhmedov ; Koksma ; Jerome . Weinberg’s no-go theorem Weinberg further shows that achieving a small cosmological constant in local, translationally invariant QFT typically requires unwanted fine tuning.
@@ III Quasi-local four-field interaction
though alternatives with similar decay yield the same qualitative result. The scale $A$ sets the range of the quasi-locality.
@@ IV Momentum space resolution of the added term
$\phi(x)=\frac{1}{V_{4}}\sum_{k}\tilde{\phi}(k)e^{ik\cdot x}\>\>,\>\>\qquad k=(\omega,\bm{k})$ (12)
@@ V Summary of the Approximations
    We replace higher time derivatives by their leading contributions in the high-$k$ regime, where spatial derivatives dominate.
@@ VI The full Lagrangian, Hamiltonian and Quartic-Oscillator Form
Additionally, combining the free Hamiltonian and the interaction term yields
@@ XI Vacuum energy computation
Substituting $k_{\rm Pl}\sim 10^{19}$ GeV and $k_{\rm cut}\sim 10^{-17}$ GeV yields a vacuum-energy density $\rho_{\rm vac}\sim 10^{-34}\,{\rm GeV}^{4}$, which is substantially reduced compared to the naive estimate and closer to the observed dark-energy scale.

### v2

@@ Abstract
In this work I present a concrete four-dimensional scalar field theory with a quasi-local, marginal (by power-counting) four-field interaction whose momentum-dependent term generates a dominant $k^{8}$ contribution to the effective Hamiltonian of each Fourier mode at large wave number $k$. When the field is confined to a finite spacetime box of side $L$, the resulting dynamics render each high-$k$ mode an effective quartic oscillator rather than a harmonic one. A semiclassical estimate of the ground-state energy then enforces an emergent UV cutoff
@@
I discuss stability, the approximations used, and generalizations to other higher-order interactions. I then compute the resulting vacuum-energy density using dimensional regularization and propose a possible way to deduce the scaling of the effective equation of state with $L$. I then include qualitative comparison with the DESI indication of mildly time-varying dark energy. The aim is to provide a field-theoretic mechanism illustrating how quasi-local interactions can generate UV/IR mixing and dynamically modify the UV cutoff scale.
@@ I Introduction
$\rho_{\rm vac}^{\rm naive}\sim\frac{1}{2}\int^{\Lambda}\!\frac{d^{3}\vec{k}}{(2\pi)^{3}}\,\omega_{k}\;\sim\;\Lambda^{4}$ (1)
with $\omega_{k}=({\vec{k}}^{2}+m^{2})^{1/2}$. Choosing $\Lambda$ near the Planck scale overshoots the observed vacuum energy by about 120 orders of magnitude. Lorentz-invariant regularization schemes such as dimensional regularization preserve $p=-\rho$ for the vacuum state and significantly reduce the result, yet still leave a discrepancy Akhmedov ; Koksma ; Jerome . Weinberg’s no-go theorem Weinberg further shows that achieving a small cosmological constant in local, translationally invariant QFT typically requires unwanted fine tuning.
@@ III Quasi-local four-field interaction
though alternatives with similar decay (such as $C(x)=\exp[-A^{4}(x.x)^{2}]\Theta(x^{2})$) yield the same qualitative result and preserve causality. The scale $A$ sets the range of the quasi-locality.
@@ IV Momentum space resolution of the added term
$\phi(x)=\frac{1}{V_{4}}\sum_{k}\tilde{\phi}(k)e^{ik\cdot x}\>\>,\>\>\qquad k=(\omega,\vec{k})$ (12)
@@ V Summary of the Approximations
    We replace higher time derivatives by their leading contributions in the high-$k$ regime, where spatial derivatives dominate. We use this to derive an approximate (and not manifestly Lorentz invariant) Hamiltonian.
@@ VI The full Lagrangian, Hamiltonian and Quartic-Oscillator Form
Additionally, combining the free Hamiltonian and the interaction term (and dropping time derivatives, as mentioned in point (2) in the previous section) yields
@@ XI Vacuum energy computation
Substituting $k_{\rm Pl}\sim 10^{19}$ GeV and $k_{\rm cut}\sim 10^{-17}$ GeV yields a vacuum-energy density $\rho_{\rm vac}\sim 10^{-34}\,{\rm GeV}^{4}$, which is substantially reduced compared to the naive estimate and closer to the observed dark-energy scale. We can also repeat the calculations in Akhmedov ; Koksma ; Jerome for all the particle species and obtain $\rho_{vac}^{TOTAL}\approx 3\times 10^{-33}GeV^{4}$, a positive number. Note, again, this $k_{cutoff}$ is very sensitive to the product $(AL)$, as described earlier in the paper.